The design of controls and instruments has a large effect on how easily people can use them. A student project investigated this effect by asking 25 right-handed students to turn a knob (with their right hands) that moved an indicator by screw action. There were two identical instruments, one with a right-hand thread (the knob turns clockwise) and the other with a left-hand thread (the knob turns counter-clockwise). The table below gives the times required (in seconds) to move the indicator a fixed distance.

Subject ID

Right-hand Thread

Left-hand

Thread

1

113

137

2

105

105

3

130

133

4

101

108

5

138

115

6

118

170

7

87

103

8

116

145

9

75

78

10

96

107

11

122

84

12

103

148

13

116

147

14

107

87

15

118

166

16

103

146

17

111

123

18

104

135

19

111

112

20

89

93

21

78

76

22

100

116

23

89

78

24

85

101

25

88

123

  1. The project hoped to show that right-handed people find right-hand threads easier to use. State the appropriate Ho and Ha about the mean time required to complete the task.
  2. Carry out a test of your hypotheses. Give the P-value, report and explain your conclusion in a brief paragraph.
Answer #1

Sabution Subyet Right -hand Thad Left hand d Right hand Thread -left hand Thread Thaead TD 13 24 137 os oS 30 133 -3 3 lo8 -7Noc to find the d a) - 333 25 d-13.32 Now to find the standand deviahion Idi-) Sd= (0+13.32)+(-3+ 13.32 ) + (=1+13.32) t1613.-13.32 t- 22.936 5 -13.32 4.58 12 t =-2.90 Now to find the degnees of faredom- |- u =P 25- df Nos to find the PValue :- fheor